OpenStudy (anonymous):

A kindergarten school has square mats in two colors, red and blue. The ratio of the perimeter of the red mat to the perimeter of the blue mat is 1 : 2. The area of the red mat is 9 square feet. What is the area of the blue mat?

OpenStudy (anonymous):

(answer choice) 9 square feet 18 square feet 81 square feet 36 square feet

OpenStudy (anonymous):

red mat:\[4\sqrt{9}=12 \]blue mat:\[4\sqrt{x}=24 \]\[\sqrt{x}=6\]\[x=36 \]